The solubility of a gas in H2O at a temperature of 25 °C and a pressure of 0.65 atm is 0.125 M. What will the solubility (in atm) of the same gas be in H2O at the same temperature at 3.7 atm?
Answer #1

The Henry's law constant "k" is different for every gas, temperature and solvent. The units on "k" depend on the units used for concentration and pressure.

The value for k is the same for the same temperature, gas and solvent. This means the concentration to pressure ratio is the same when pressures change. The following equation can be used to relate pressure and concentration changes for solutions at the same temperature. The initial condition has concentration C1 and gas partial pressure P1. The second condition has concentration C2 and gas partial pressure P2.

Henrys-law-ratio,

0.125/0.65= C2/3.7

C2=0.125x3.7/0.65=0.7115 M